A population consists of the following measurement 3,8,14,17,and 58 consider all the sample of size 3 taken from this population without replacement



(1)List the sample of the mean and media respectively

(2) calculate the mean of the sample distribution of the media and the means


1
Expert's answer
2021-04-15T06:50:05-0400


(1) Mean

"\bar{X} = \frac{3+8+14+17+58}{5} = 20"

Median = 14

(2) n = 60

"Mean_{sd} = \frac{\sum Mean_i}{60} = 20.07"

Median = 24 (by Excel function)